# Cost-Optimal Microservices Deployment with Cluster Autoscaling and Spot Pricing

### v1

@@ III-E Cluster Auto Scaling
Based on the current pod count and other configurations, such as spot termination details, the optimization engine will recommend a new optimal node combination to the node allocator and resources will then be attached to or detached from the cluster.
Following flowchart 4 illustrate the high level workflow of the elastic scaler.
Figure 4: Elastic Scaler
@@ IV-A1 Hardware setup
The Kubernetes manager node was provisioned on an AWS EC2 instance of type t3.medium. This instance type was selected due to its balanced compute, memory, and networking resources, aligning with the requirements of managing the Kubernetes cluster efficiently.
@@
A t3.medium EC2 instance was designated as the Kubernetes worker node. This instance type provides a suitable balance of resources to support the execution of containerized workloads within the Kubernetes cluster while remaining cost-effective.
@@
The SpotKube services, responsible for managing AWS spot instances within the Kubernetes cluster, were deployed on a t2.micro EC2 instance. Despite its smaller resource footprint compared to other instance types, the t2.micro instance was deemed sufficient for hosting the SpotKube services without compromising their functionality.
@@ IV-A2 Microservice applications
HTTP-based microservices are used for the experiment as HTTP-based microservices facilitate the evaluation of the impact of network latency and bandwidth on the performance of the microservices. Response times and network traffic were measured to identify potential bottlenecks in the system and assess the effectiveness of SpotKube in optimizing resource allocation to alleviate these issues. The following four microservices are selected from popular open-source projects for the experiments.
  • Gateway service (Network heavy application) - Work as a proxy and forward the traffic to internal services.
  • Auth service (Database access) - Query the MySQL database and validate the username and password.
  • Guest-book Go service (Web access) - Dynamically generate HTML content based on the users’ input.
  • Prime-app service (Compute intensive) - Calculate all prime numbers up to a given number.
Additionally, each microservice should have a load-testing script and user-defined Service Level Objectives.
@@ IV-A3 Load testing
In order to conduct the load test, Spotkube utilized the Locust test script provided by the user. Locust linearly increases the number of users spawned per second. Prometheus collected CPU metrics, and Locust collected test statistics and saved them into a CSV file for analysis.
@@ IV-A4 Deployment Strategy
Resource limits were imposed on each container to maintain simplicity, and Minikube was selected as the Kubernetes cluster for the experiment. Two deployment strategies were utilized to compare the effectiveness of the proposed approach.
In the first strategy, the Kubernetes horizontal pod autoscalar(HPA) was used to deploy the microservices along with the Cluster Autoscaler(CA) with AWS Autoscaling group, and subsequently, the applications were subjected to load testing.
In the second strategy, both HPA and SpotKube were used to deploy the microservices, and the applications were subjected to load testing.
@@ IV-B1 Analytical Model Validation
Assuming that the user-defined SLO for the Auth service is to support a minimum of 50 requests per second, the load test results for this service that was deployed using the first strategy which is enabling the HPA and AWS CA are depicted in Figure 5. However, it can be observed that this service encountered some failures before reaching the specified SLO.
On the other hand, Figure 6 displays the load test results for the same Auth service that was deployed using Spotkube and HPA, where it was able to achieve the user-defined objectives.
Figure 5: Total requests per second for default configuration
Figure 6: Total requests per second for SpotKube configuration
Table I & Table II show a comparison between the two deployment strategies. Each service user has specified the minimum number of requests that should support per second. These data indicate that using Spotkube to deploy microservices is more effective than deploying microservices in an ad-hoc manner.
Service Load testing SLO (RPS) Default
Number
of users
Spawn
rate
RPS when
observe failures
Number of
@@
Auth 1000 10 45 43,2 1.6
Gateway 1000 10 50 46.5 6.5
Guestbook 1000 10 50 43.3 2.8
Prime-app 500 2 60 51.2 0
TABLE I: HPA and CA - Default Behavior
Service Load testing SLO (RPS) Spotkube
Number
of users
Spawn
rate
RPS when
observe failures
Number of
@@
Auth 1000 10 45 49.4 1.5
Gateway 1000 10 50 55.1 1.9
Guestbook 1000 10 50 54.1 1.6
Prime-app 500 2 60 72.5 0
TABLE II: HPA and SpotKube
@@ IV-B2 Spot Predictor Validation
SpotKube relies on the predicted spot prices when requesting spot instances from the AWS spot market. Hence, the accuracy of the prediction is crucial to mitigate unnecessary terminations of spot instances. We evaluate time series model accuracy for two instance types, and the results show that the prediction was accurate enough to bid spot instances from the spot market (Figures: 7, 8).
Figure 7: Root Mean Square Error: 0.00341179
Figure 8: Root Mean Square Error: 0.00150967
In summary, forecasting models are able to predict the spot price properly and these values can use to request spot instances from the spot market. Root Mean Square Error proves the effectiveness of each prediction.
@@ IV-B3 Optimization Algorithms Comparison
We conducted an experiment to compare the performance of three optimization algorithms: Bruteforce, Greedy approach, and Genetic algorithm-based Pareto optimization. The objective of the experiment was to evaluate the algorithms based on three criteria: execution time, total cost, and the number of nodes.
Throughout the experiment, we analyzed the algorithms’ performance with varying numbers of pods. By considering different numbers of pods, we aimed to understand how the algorithms’ efficiency and effectiveness change in response to the scale of the problem.
  • pymoo: Pareto optimization
  • greedy: Greedy Approach
  • bruteforce: Bruteforce Approach
Time Comparison
According to the below graph,
  • pymoo - Regardless of the pod count, execution time is almost the same throughout the experiment
  • greedy - Execution time is less compared to other algorithms.
  • bruteforce - Exponentially grows with the pod count
Figure 9: Time Comparison
Outcome: Greedy approach performs better. Pymoo also comparatively performing well
Hourly Cost (\$) Comparison
According to the below graph,
  • pymoo - Returns the lowest cost
  • greedy - Cost is not optimal
  • bruteforce - Gives the optimal cost. But when the pod count increases, the cost retuned by the brute force approach is greater than the cost returned by the Pareto-based implementation. This is due to the constraint that has been made during the implementation of the brute force approach to make it more efficient.
Figure 10: Total Hourly Cost Comparison
Outcome: Pymoo performs well by identifying the most cost-optimal node combination
Node Count Comparison
According to the below graph,
  • pymoo - Returns the maximum number of nodes
  • greedy - Returns a higher number of nodes
  • bruteforce - A lower number of node count is returned compared to other approaches
Figure 11: Node Count Comparison
Outcome: Pymoo performs well by identifying the maximum possible node count
Summary
The primary goal of the optimization algorithm is to identify the best trade-off between cost and performance and output the optimal set of node combinations. Cost is calculated using the defined cost model, while the performance is guaranteed by deploying the pods on many nodes. In other words, distributing pods as much as possible. This will ensure that the availability of the system is maintained even in a sudden termination of nodes.
According to the above experiment, Pareto optimization was able to output the best set of nodes where the cost is minimal, and the node count is maximal.
Prime app 2 6 1 (t3.medium) 2 (t3.medium, t3.medium) 8.1
Prime app 2 10 1 (t3.medium) 3 (t3.medium, t3.medium, m6g.medium) 7.2
Gateway 8 4 2 (t3.medium, t3.medium) 1 (t3.medium) 7.7
TABLE III: Elastic Scaler Efficiency Validation
@@ IV-B4 Elastic Scaler Validation
Elastic Scaler Efficiency Validation
Elastic Scaler and the Optimization Engine decision were taken within 30 seconds while node allocation process took around 6-7 minutes. In comparison, the AWS Cluster Autoscaler with AWS Autoscaling group took less than 30 seconds to make scaling decisions, while the cloud provider required 2 to 8 minutes to create compute resources [20] [21]. Consequently, further research is necessary to decrease the cluster autoscaling time and enhance the performance of node allocation process.
@@
To assess the cost-effectiveness of our approach, we conducted a comparison between our SpotKube solution and Amazon EKS utilizing spot instances and the Cluster Autoscaler (CA). The Horizontal Pod Autoscaler (HPA) was enabled in both cases. For the comparison, we initially deployed two applications, Prime App and Gateway Service, utilizing the elastic scaler and other SpotKube services. We then sent requests to both applications according to a known pattern for 24 hours. Similarly, we performed the same deployment using EKS and CA, followed by the same request-sending process for 24 hours. We restricted the resource limit of each pod to 0.5 vCPU and 1 GiB of memory. For the node group, we chose the t3.medium, c6a.large, t4g.large, and c6g.xlarge spot instance types.
After the experiment, we calculated that the average pod count per hour was around 50 for both deployments. The resulting average spot instance count and associated cost are presented in the table below, with the left side indicating the cost for the Amazon EKS cluster and the right side representing the cost for the SpotKube cluster.
@@ V DISCUSSION
In this discussion, we critically analyze and delve into the effectiveness of SpotKube, a framework designed to address these challenges through innovative approaches such as application characterization, elastic cluster autoscaling, and integration of spot pricing models.
@@ V-A Effectiveness of Application Characterization
Application characterization is responsible for determining the initial number of pods (replicas) that need to be deployed without violating the user-defined SLOs. Due to the dynamic nature of microservice applications, SpotKube is only considered for CPU-intensive applications, and for the deployment, homogeneous pods are considered. As described in the methodology section, each microservice is subjected to a load test, and results were analyzed to determine the required pod count without violating SLOs. The data we collected through the load test consist of CPU usage and the failure rate per given load (RPS). Through experiments, we identified that by analysing the failure rates and CPU usage data related to the load test, we can determine the maximum RPS value that can be handled by a pod. Then, by mapping this RPS value with the SLO (Minimum RPS value that a microservice should handle), we were able to determine the initial pod count that needed to be deployed. Experimental results show that the default behaviour of Kubernetes with the HPA encountered some failures during pod scaling, while SpotKube is able to satisfy the user-defined SLOs without any failure. Although this can be thought of as an over-provisioning mechanism, due to the usage of spot instances, cost has been reduced significantly, and the application performance is maintained without any failures.
@@ V-B Elastic Cluster Autoscaling through Pareto Optimization
The elastic resource allocation mechanism within the SpotKube dynamically adjusts the deployment of applications based on microservices architecture, taking into consideration the current spot market prices and the demand of the application. The elastic cluster autoscaler employs the default Horizontal Pod Autoscaler (HPA) for the autoscaling of pods. This process is triggered once the total CPU utilization of the cluster surpasses an 80% threshold or falls below a 30% threshold, at which point the optimization engine is invoked to perform the necessary resource allocation or deallocation within the cluster. The determination of these specific threshold values, 80% and 30%, was informed by an extensive review of existing literature as well as iterative trials and errors, to maintain the required performance of the application.
A significant contribution of the SpotKube framework is the utilization of a genetic algorithm for Pareto optimization, which facilitates an exploration of the cost-performance trade-offs to identify optimal deployment configurations. The experimental evaluation of this approach highlights its efficacy in identifying efficient solutions, with a particular focus on the algorithm’s execution time. According to the Time Comparison graph, Pareto optimization consistently determines the most cost-effective and performance-optimal spot instances that should be attached to the cluster within a nearly uniform duration across various deployment scenarios. This aspect of the optimization algorithm, which strategically determines the most cost-effective set of spot instances while maximizing the count of instances to distribute the workload broadly, represents a proactive strategy to mitigate the risk associated with the abrupt termination of spot instances.
The capability of the cluster autoscaler to dynamically scale clusters is evident, yet it is not without its limitations. Specifically, in the context of Amazon EKS, the autoscaler requires a minimum of 2 minutes to integrate a new node into the cluster, with the process potentially extending up to 8 minutes. Conversely, experimental observations have indicated that the Spotkube Elastic Scaler averages a scaling-up period of about 6-8 minutes. These experiments were conducted using an Ubuntu AMI, onto which all necessary software packages were installed via Ansible. This approach significantly leads to the extended duration required for scaling operations. We assume that employing a custom AMI, pre-equipped with all requisite packages, could notably accelerate the scaling process.
@@ V-C Integration of Spot Pricing and Cost -Effectiveness of SpotKube
Abrupt termination of spot instances is a significant challenge that can cause substantial downtime in the cluster, directly impacting system availability. To mitigate this, SpotKube employs a mix of novel and traditional techniques. One strategy is to strategically request spot instances at a bid higher than the forecasted spot price, yet still below the cost of on-demand instances. Another technique involves allocating a mix of spot instances from different instance families and availability zones. Furthermore, our optimization strategy focuses on minimizing costs while maximizing the total node count. By maximizing node count, SpotKube distributes the workload as much as possible, thereby maintaining system availability even with the sudden termination of a node. Additionally, SpotKube utilizes AWS’s abrupt termination signals to handle node termination and pod rescheduling gracefully. This is managed by the elastic scaler and optimization engine.
@@
VI CONCLUSIONS and FUTURE WORKS
@@
As future work, we plan to enhance SpotKube’s capabilities to support for heterogeneous pod deployment. which will enable SpotKube to optimize the deployment of applications that use different hardware configurations, enabling better resource utilization and cost savings. We aim to enhance the performance of Elastic Scalar by minimizing the time it takes for scaling operations.
In summary, this research demonstrates the potential of leveraging spot pricing in public cloud environments to reduce costs through cluster auto scaling while maintaining application performance, and future works aim to improve and expand SpotKube’s capabilities to further optimize microservices deployment in cost-optimized manner.

### v2

@@ III-E Dynamic Scaling of the Cluster
Based on the current pod count and spot termination details, the optimization engine will recommend a new optimal node combination and resources will then be attached to or detached from the cluster accordingly.
@@ IV-A1 Hardware setup
The Kubernetes manager node was provisioned on an AWS EC2 instance of type t3.medium which aligns with the requirements of managing the Kubernetes cluster efficiently.
@@
A t3.medium EC2 instance was designated as the Kubernetes worker node.
@@
The SpotKube services, responsible for managing AWS spot instances within the Kubernetes cluster, were deployed on a t2.micro EC2 instance.
@@ IV-A2 Microservice applications
The following four HTTP-based microservices are selected from popular open-source projects for the experiments.
  • Gateway service (Network heavy application) [20]
  • Auth service (Database access) [20]
  • Guest-book Go service (Web access) [21]
  • Prime-app service (Compute intensive)
@@ IV-A3 Load testing
The microservices used in this experiment operate asynchronously. For synchronous services, mock implementations of dependencies would be necessary before load testing, as suggested by Jindal et al. [2]. To conduct load testing, SpotKube utilizes Locust, which linearly increases the number of users spawned per second, while Prometheus collects CPU metrics.
@@ IV-A4 Deployment Strategy
Resource limits were imposed on each container for simplicity, and Minikube was selected as the K8s cluster for the experiment. Two deployment strategies were employed followed by load testing of the applications to evaluate the proposed approach:
  • Default - Horizontal Pod Autoscaler (HPA) and Cluster Autoscaler (CA) with AWS Autoscaling Group
  • SpotKube - HPA and SpotKube
@@ IV-B1 Analytical Model Validation
For this evaluation, we set a user-defined Service Level Objective (SLO) for all applications, targeting a minimum support of 50 requests per second.
Figure 4 presents the load test results for the Auth service, which was deployed using both default and SpotKube strategies. The results indicate that the default deployment setup experienced some failures before meeting the specified SLO. In contrast, the SpotKube-based deployment successfully achieved the user-defined objectives.
Figure 4: SpotKube vs Default Behavior of Auth Service
Table I compares two deployment strategies, highlighting the RPS at which failures occur with the Service Level Objective (SLO). The results indicate that deploying microservices using SpotKube is more effective than an ad-hoc deployment approach.
Service Spawn Rate SLO (RPS) Default Spotkube
RPS
@@
RPS
@@
Auth 10 50 46.1 4.3 55.1 1.9
Gateway 10 50 46.5 6.5 54.1 1.9
Guestbook 10 50 43.3 2.8 54.1 1.6
Prime-app 2 50 51.2 0 62.5 0
TABLE I: Analytical Model Validation
@@ IV-B2 Spot Predictor Validation
Figure 5: Spot Price Prediction
SpotKube relies on accurate spot price forecasts to request spot instances from the AWS spot market, helping to minimize unnecessary instance terminations. As shown in Figure 5, we evaluated our Prophet-based spot price forecasting model on two instance types. The model’s low Root Mean Square Error (RMSE) values indicate high forecasting accuracy, supporting reliable spot instance bidding. Furthermore, the narrow 95% confidence intervals reinforce the robustness of the approach, confirming their reliability for spot price prediction.
@@ IV-B3 Optimization Algorithms Comparison
We experimented to compare the performance of three optimization algorithms: Bruteforce, Greedy approach, and NSGA-II-based Pareto optimization. The evaluation focused on three criteria: execution time, total cost, and the number of nodes. The algorithms’ performance was analyzed with varying numbers of pods to understand their efficiency and effectiveness relative to the problem’s scale.
The primary objective was to identify the optimal trade-off between cost and performance by determining the best set of node combinations, where cost is calculated using a defined cost model, and performance is ensured by distributing pods across multiple nodes to maintain system availability even during sudden spot instance terminations.
According to Figure 6, results indicated that Pareto optimization consistently produced the best set of nodes in a uniform time frame, regardless of the pod count, achieving minimal cost and maximal node distribution.
(a) Time(s)
(b) Instance Count
(c) Hourly Cost(\$)
Figure 6: Optimization Algorithm Comparison
@@ IV-B4 Elastic Scaler Validation
To assess the cost-effectiveness of our Elastic Scaler and SpotKube solution, we compared it to an Amazon EKS cluster [22] based on spot instances, with both the Cluster Autoscaler (CA) and Horizontal Pod Autoscaler (HPA) enabled. The experiment involved deploying two applications, the Prime App and the Gateway Service, with requests sent to both applications according to a predetermined pattern over 24 hours.
Resource limits for each pod were set to 0.5 vCPU and 1 GiB of memory, with the node group comprising t3.medium, c6a.large, t4g.large, and c6g.xlarge spot instance types. Post-experiment analysis revealed that the average pod count per hour was approximately 50 for both deployments. The average spot instance count and the associated costs were documented, with comparative costs for the Amazon EKS cluster and the SpotKube cluster presented in Table II.
@@
Prime app 2 6 1 (t3.medium) 2 (t3.medium, t3.medium) 8.1
Prime app 2 10 1 (t3.medium) 3 (t3.medium, t3.medium, m6g.medium) 7.2
Gateway 8 4 2 (t3.medium, t3.medium) 1 (t3.medium) 7.7
TABLE III: Elastic Scaler Efficiency Validation
Elastic Scaler Efficiency Validation
As depicted in Table III Elastic Scaler and the Optimization Engine decision were taken within 30 seconds while the node allocation process took around 6-7 minutes. In comparison, the AWS Cluster Autoscaler with AWS Autoscaling group took less than 30 seconds to make scaling decisions, while the cloud provider required 2 to 8 minutes to create compute resources [23] [24]. Consequently, further research is necessary to decrease the cluster autoscaling time and enhance the performance of the node allocation process.
@@ V DISCUSSION
In this discussion, we critically analyze the effectiveness of SpotKube, a framework designed to tackle challenges related to cloud cost optimization. It employs innovative approaches such as application characterization, elastic cluster autoscaling, and the integration of spot pricing models.
@@ V-A Effectiveness of Application Characterization
Application characterization is responsible for determining the initial number of pods (replicas) that need to be deployed without violating the user-defined SLOs. Due to the dynamic nature of microservice applications, SpotKube is only considered for CPU-intensive applications, and for deployment, homogeneous pods are considered. As described in the methodology section, each microservice is subjected to a load test, and results were analyzed to determine the required pod count without violating SLOs. The data we collected through the load test consists of CPU usage and the failure rate per given load (RPS). Through experiments, we identified that by analyzing the failure rates and CPU usage data related to the load test, we can determine the maximum RPS value that can be handled by a pod. Then, by mapping this RPS value with the SLO (Minimum RPS value that a microservice should handle), we were able to determine the initial pod count that needed to be deployed. Experimental results show that the default behavior of Kubernetes with the HPA encountered some failures during pod scaling, while SpotKube is able to satisfy the user-defined SLOs without any failure. Although this can be thought of as an over-provisioning mechanism, due to the usage of spot instances, the cost has been reduced significantly, and the application performance is maintained without any failures.
@@ V-B Elastic Cluster Autoscaling through Pareto Optimization
The resource allocation mechanism within SpotKube, dynamically adjusts the deployment of microservice applications, taking into consideration the current spot market prices and the demand of the application. The elastic scaler employs the default Horizontal Pod Autoscaler (HPA) for the autoscaling of pods. This process is triggered once the total CPU utilization of the cluster surpasses an 80% threshold or falls below a 30% threshold, at which point the optimization engine is invoked to perform the necessary resource allocation or deallocation within the cluster. The determination of these specific threshold values, 80% and 30%, was informed by an extensive review of existing literature as well as iterative trials and errors, to maintain the required performance of the application.
A significant contribution of SpotKube is the utilization of a genetic algorithm for Pareto optimization, which facilitates an exploration of the cost-performance trade-offs to identify optimal deployment configurations. The experimental evaluation of this approach highlights its efficacy in identifying efficient solutions, with a particular focus on the algorithm’s execution time. According to the Time Comparison graph, Pareto optimization consistently determines the most cost-effective and performance-optimal spot instances that should be attached to the cluster within a nearly uniform duration across various deployment scenarios. This aspect of the optimization algorithm, which strategically determines the most cost-effective set of spot instances while maximizing the count of instances to distribute the workload broadly, represents a proactive strategy to mitigate the risk associated with the abrupt termination of spot instances.
The capability of the elastic scaler to dynamically scale clusters is evident, yet it is not without its limitations. Specifically, in the context of Amazon EKS, the autoscaler requires a minimum of 2 minutes to integrate a new node into the cluster, with the process potentially extending up to 8 minutes. Conversely, experimental observations have indicated that the Spotkube Elastic Scaler averages a scaling-up period of about 6-8 minutes. These experiments were conducted using an Ubuntu AMI, onto which all necessary software packages were installed via Ansible [25]. This approach significantly leads to the extended duration required for scaling operations. We assume that employing a custom AMI, pre-equipped with all requisite packages, could notably accelerate the scaling process.
@@ V-C Integration of Spot Pricing and Cost -Effectiveness of SpotKube
Abrupt termination of spot instances is a significant challenge that can cause substantial downtime in the cluster, directly impacting system availability. To mitigate this, SpotKube employs a mix of novel and traditional techniques. One strategy is to strategically request spot instances at a bid higher than the forecasted spot price, yet still below the cost of on-demand instances. Another technique involves allocating a mix of spot instances from different instance families and availability zones. Furthermore, our optimization strategy focuses on minimizing costs while maximizing the total node count. By maximizing node count, SpotKube distributes the workload as much as possible, thereby maintaining system availability even with the sudden termination of a node. Additionally, SpotKube utilizes AWS’s abrupt termination signals to handle node termination and pod rescheduling gracefully.
@@
VI CONCLUSION and FUTURE WORKS
@@
As future work, we plan to enhance SpotKube’s capabilities to support heterogeneous pod deployment. This will enable SpotKube to optimize the deployment of applications that use different hardware configurations, enabling better resource utilization and cost savings. We aim to enhance the performance of Elastic Scalar by minimizing the time it takes for scaling operations.
In summary, this research demonstrates the potential of leveraging spot pricing in public cloud environments to reduce costs through cluster auto-scaling while maintaining application performance, and future works aim to improve and expand SpotKube’s capabilities to further optimize microservices deployment in a cost-optimized manner.